Elliptic Operators on Planar Graphs: Unique Continuation and Curvature

Steffen Klassert TU Chemnitz, Germany

This talk is concerned with elliptic operators on plane tesselations. We discuss that such an operator does not admit a compactly supported eigenfunction, if the combinatorial curvature of the tesselation is nonpositive.